OpenStudy (oswaldmurphy):

Given the following functions f(x) and g(x), solve f over g(−4) and select the correct answer below: f(x) = 4x − 4 g(x) = x − 1

OpenStudy (oswaldmurphy):

−20 −4 4 one fourth @Jadeishere

OpenStudy (jadeishere):

Eeeeks, I hate these, I don't know if I'd do it right @TheSmartOne

OpenStudy (slenderman):

random guess,....its seven

OpenStudy (faiqraees):

Okay so f(x) over g(x) imply f(x)/g(x)

OpenStudy (faiqraees):

so f(x)/g(x) = (4x-4)/(x-1) =4(x-1)/(x-1) =4

OpenStudy (faiqraees):

Understood or need an explanation?

OpenStudy (sila1453):

g(x)= x-1 put x=-4 in the above equation f(x)/g(x)= 4x-4 / x-1 =4(x-1)/x-1 = 4 f(x)/g(-4)= 4 is your correct answer

OpenStudy (slenderman):

so its definatley 7 ok thx, hahah jk hes right

OpenStudy (oswaldmurphy):

@FaiqRaees can you explain a little bit? More on the formula?

OpenStudy (faiqraees):

Okay so lets just modify the question a bit y=4x-4 z=x-1 find y/z when x=-4 =y/z = (y=4x-4)/(z=x-1) [This type of substitution is just done to explain clearly] = 4x-4 / x-1 Now simplify the above expression.

OpenStudy (oswaldmurphy):

Hm,... Is it 4?